P^3x+p^2y-p^2x-py=0, where p=dy/dx


1
Expert's answer
2021-07-06T05:22:34-0400

Given equation is: "p^3x+p^2y-p^2x-py=0"


"p^2x(p-1) + py(p-1) = 0"

"p(p-1)(px+y) = 0"


Solving it we get, "p=0,p-1 = 0, px+y=0"


Then, for "p = 0, \implies \frac{dy}{dx} = 0"

Integrating it we get, "y = c_1" (1)

where c1 is constant.


For p-1=0,

"\frac{dy}{dx} -1= 0"

"\int {dy} =\int dx"

"y = x + c_2" (2)



For px+y = 0,

"x\frac{dy}{dx} + y= 0"


"x\frac{dy}{dx} = -y"


"\frac{dy}{y} = -\frac{dx}{x}"

integrating both sides,

"\ln y = -\ln x + \ln c_3"

"y = \frac{c_3}{x}" (3)



Equations (1), (2), and (3) are the solutions of the differential equation.